OpenStudy (anonymous):

Want a medal? What is the coefficient of the x^9y -term in the binomial expansion of (2y + 4x^3)4? 4 32 128 512

OpenStudy (anonymous):

512\[256 x^{12}+512 x^9 y+384 x^6 y^2+128 x^3 y^3+16 y^4 \]

OpenStudy (anonymous):

@robtobey Um, I'm confused on how you got your answer. Could you explain, please?

OpenStudy (anonymous):

\[\left(4 x^3+2 y\right) \left(4 x^3+2 y\right) \left(4 x^3+2 y\right) \left(4 x^3+2 y\right) \]

OpenStudy (anonymous):

If you running on a Mac, use the Firefox browser, for viewing formulas. Chrome is not rendering correctly on my machine.

OpenStudy (anonymous):

okay then what

OpenStudy (anonymous):

\[\left(4 x^3+2 y\right) \left(4 x^3+2 y\right)=16 x^6+16 x^3 y+4 y^2 \]\[\left(16 x^6+16 x^3 y+4 y^2\right) \left(16 x^6+16 x^3 y+4 y^2\right)= \]\[256 x^{12}+512 x^9 y+384 x^6 y^2+128 x^3 y^3+16 y^4 \]